121314gifted:

The formula: d = rt is used to find the distance. Write a RELATED EQUATION to solve for the variable, t.

Shadow:

What they are basically asking you to do is isolate t, to write an equation that basically has t alone \[t = ?\]

Shadow:

@121314gifted Do you know where to start with that?

121314gifted:

Not really, see sometimes I'm good in Math and other times I'm not, its complicated

Shadow:

Say they gave me the formula x = q(p/a), and they want me to find the related equation for a, here's how: \[x = q (\frac{ p }{ a })\] Divide both sides by q \[\frac{ x }{ q } = \frac{ p }{ a }\] Multiply both sides by 1/p \[\frac{ x }{ qp } = a\]

Shadow:

You just have to use Algebra to get that variable alone basically.

121314gifted:

Thxs Shadow your really good at this!

121314gifted:

I appreciate the help

Shadow:

No problem